OpenStudy (anonymous):

What is 27^(2x) = 9^(x-3)

OpenStudy (caishaax3):

do you need x?

OpenStudy (anonymous):

Yes. Sorry I forgot to write that.

OpenStudy (caishaax3):

x=-3/2

OpenStudy (jdoe0001):

what's \(3^2=?\) and, say what's \(3^3=?\)

OpenStudy (anonymous):

Can you explain how to get that?

OpenStudy (jdoe0001):

\(\large { {\color{brown}{ 27}}^{2x} = {\color{brown}{ 9}}^{x-3}\qquad \begin{cases} 3^2\to {\color{brown}{ 9}}\\ 3^3\to {\color{brown}{ 27}} \end{cases}\qquad thus\implies (3^3)^{2x}=(3^2)^{x-3} }\) see how to get "x" now?

OpenStudy (anonymous):

Um...not exactly sure....Am I suppose to combine the 2x and x-3?

OpenStudy (jdoe0001):

\(\large { {\color{brown}{ 27}}^{2x} = {\color{brown}{ 9}}^{x-3}\qquad \begin{cases} 3^2\to {\color{brown}{ 9}}\\ 3^3\to {\color{brown}{ 27}} \end{cases}\qquad thus\implies (3^3)^{2x}=(3^2)^{x-3} \\ \quad \\ 3^{3\cdot (2x)}=3^{2\cdot (x-3)}\impliedby \begin{array}{llll} \textit{same base, thus}\\\ \textit{exponents must also equal each other} \end{array}\\ \quad \\ 3(2x)=2(x-3) }\)

OpenStudy (jdoe0001):

\(\large a^{whatever} = a^{whatever}\) means that whatever = whatever

OpenStudy (anonymous):

Ohh! So then: 6x=2x-6 4x=-6 x=-2/3

OpenStudy (anonymous):

Is that right?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

hmmm actually wait

OpenStudy (anonymous):

Thanks!

OpenStudy (jdoe0001):

\(\bf 6x=2x-6\implies 4x=-6\implies x=-\cfrac{\cancel{6}}{\cancel{4}}\implies x=-\cfrac{3}{2}\)

OpenStudy (anonymous):

Oh whoops! Thanks for catching that!